A rectangle is drawn so the width is 9 inches longer than the height. If the rectangle's diagonal measurement is 45 inches, find the height.
Give your answer rounded to 1 decimal place.

Answer by Alan3354(69443)

A rectangle is drawn so the width is 9 inches longer than the height. If the rectangle's diagonal measurement is 45 inches, find the height.
----------
w = h + 9
w^2 + h^2 = 45^2
=========================
Give your answer rounded to 1 decimal place.
Couldn't you do the rounding?